OpenStudy (anonymous):

In 1980 Median family income was about $19,000 and in 2000 it was about $43,000. Find the slope of the line passing through points (1980,19000) and (2000,43000). Interpret the slope as a rate change. What will the median family income be in 2007?

OpenStudy (anonymous):

y2-y1 over x2-x1 is slope

OpenStudy (anonymous):

Can you please help me solve this?

OpenStudy (anonymous):

okay so do (43000-19000)/(2000-1980) to get slope use y-y1=m(x-x1) to find the equation (solve for y) plug in your new x

OpenStudy (anonymous):

m=24000/20 =1200

OpenStudy (anonymous):

The family income increases by $1200 between 1980 and 2000. Right?

OpenStudy (anonymous):

each year it increases that much

OpenStudy (anonymous):

So the change is..... the family income is increasing at the rate of $1200 per year

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

What will be the family median income in 2007?

OpenStudy (anonymous):

pick a set of coordinates to be your (x1,y1) plug into y-y1=m(x-x1) to get something in the form of y=mx+b plug in 2007 for the x in your new equation

OpenStudy (anonymous):

If the trend continues, estimate the family median in 2007. Do I use the same equation? Instead of using 2000-1980. use 2007-1980

OpenStudy (anonymous):

standby

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

http://imgur.com/OzsZu

OpenStudy (anonymous):

solve the bottom

OpenStudy (anonymous):

the answer I got was 51,400

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Find the slope, if it exists, of the line containing the pair of points. (-15,-4) and (-17,-12) The slope is? The slope is undefined? Please help me solve this

OpenStudy (anonymous):

same thing as before

OpenStudy (anonymous):

m=-12-(-4)/-17-(-15) =16/32

OpenStudy (anonymous):

m=2/4

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

the slope is 2/4

OpenStudy (anonymous):

Am I correct?